# Optimized quantum sensor networks for ultralight dark matter detection

### v1

@@ Abstract
In Fig. 5(b), we further investigate the QB as a function of $L_{1}$. Different from four-qubit case, here, the best performance is obtained when ($L_{1}=2$). This result suggests that even for larger systems, a shallow and well-optimized circuit is sufficient to achieve high-precision estimation, which is advantageous for practical implementation on near-term quantum devices.
@@
Figure 5(c) shows how the CB varies with $L_{2}$. As $L_{2}$ increases, the CB improves and approaches the QB. However, it does not reach the QB due to limitations in the measurement strategy. In this analysis, we evaluate up to $L_{2}=3$. Remarkably, the precision in this case is better than that of the four-qubit case.
Figure 5: QBs for a nine-qubit system ($N=9$). (a) Comparison of various ansatze, following the designs shown in Fig. 1(c). Among them, F9 achieves the best QB. Compared to the four-qubit case, the QB is improved, consistent with the expectation that precision increases with the number of qubits. (b) Dependence of the QB on the number of layers $L_{1}$ for F9 configuration. The best performance is achieved with a shallow circuit at $L_{1}=2$, highlighting the efficiency and practicality of the approach for larger quantum networks. (c) CB as a function of the number of layers $L_{2}$. Increasing $L_{2}$ improves the CB, bringing it closer to the QB.
@@
Noisy effect. We analyze the efficiency of a quantum sensor network under local dephasing noise. Each sensor node (qubit) undergoes dephasing described by the Kraus operators $K_{0}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\lambda}\end{pmatrix},K_{1}=\begin{pmatrix}0&0\\ 0&\sqrt{\lambda}\end{pmatrix}$, where $\lambda\in[0,1]$ characterizes the dephasing strength. The resulting channel acts as
$\mathcal{E}(\rho)=K_{0}\rho K_{0}^{\dagger}+K_{1}\rho K_{1}^{\dagger}.$ (15)
@@
The numerical results for the QB under varying dephasing strength $\lambda$ are shown in Fig. 6 for two configurations, F4 (a) and F9 (b). No re-optimization is performed for each noise level; instead, the optimal configuration setup at $\lambda=0$ is used throughout. Remarkably, the QBs slightly increase even as the noise strength rises to $\lambda=0.9$. Specifically, the difference between the QB at $\lambda=0.9$ and the noiseless case is $\Delta=1.68\times 10^{-5}$ for F4 and $\Delta=3.73\times 10^{-5}$ for F9. This demonstrates that the optimal sensor networks remain robust under dephasing, with only minimal degradation in sensitivity even without re-optimization.
Figure 6: QBs under dephasing noise for (a) F4 and (b) F9 configurations. QBs exhibit a slight increase with increasing dephasing strength $\lambda$, indicating a modest reduction in sensitivity. However, even at high noise levels ($\lambda=0.9$), the increase remains small: $\Delta=1.68\times 10^{-5}$ for (a) and $\Delta=3.73\times 10^{-5}$ for (b).
@@
Conclusion. We proposed and demonstrated a network-based quantum sensing strategy for enhancing DM detection. By structuring superconducting qubits into networks with specific symmetries and optimizing their connections using VQM, we achieved substantial improvements in sensing precision compared to conventional GHZ-based protocols. Both the QB and CB were systematically minimized, and Bayesian estimation further validated the enhanced sensitivity of the optimized networks. Importantly, the optimal configurations maintained shallow circuit depths, making them feasible for implementation on NISQ devices. Our findings emphasize the significant impact of network design on quantum metrology performance and suggest that well-structured quantum networks can serve as powerful platforms not only for DM searches but also for a wide range of high-precision sensing applications. Future works may extend these results by exploring larger networks, adaptive strategies, and integration with error mitigation techniques to further push the limits of quantum-enhanced detection.

### v2

@@ Abstract
In Fig. 5(b), we investigate the QB as a function of $L_{1}$. Different from four-qubit case, here, the best performance is obtained when ($L_{1}=2$). This result suggests that even for larger systems, a shallow and well-optimized circuit is sufficient to achieve high-precision estimation.
Figure 4: (a) Posterior versus phase $\delta$ for different $\nu$; the peak approaches the true value as $\nu$ grows. (b) Bias $\bar{\delta}-\delta_{\rm true}$ converges to zero with $\nu$. (c) Variance Var($\delta$) versus $\nu$, compared with ${\rm CB}/\nu$ and ${\rm QB}/\nu$, showing agreement and optimal scaling.
@@
Figure 5(c) shows the CB versus $L_{2}$. As $L_{2}$ increases, the CB improves and approaches the QB. However, it does not reach the QB due to limitations in the measurement strategy. In this analysis, we evaluate up to $L_{2}=3$. Remarkably, the precision in this case is better than that of the four-qubit case.
@@
Noisy effect. We analyze the efficiency of a quantum sensor network under local dephasing noise. Each qubit undergoes dephasing described by the Kraus operators $K_{0}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\lambda}\end{pmatrix},K_{1}=\begin{pmatrix}0&0\\ 0&\sqrt{\lambda}\end{pmatrix}$, where $\lambda\in[0,1]$ characterizes the dephasing strength. The channel acts as $\mathcal{E}(\rho)=K_{0}\rho K_{0}^{\dagger}+K_{1}\rho K_{1}^{\dagger}.$
@@
The numerical results for the QB versus $\lambda$ are shown in Fig. 6 for F4 (a) and F9 (b). No re-optimization is performed for each noise level; instead, the optimal configuration setup at $\lambda=0$ is used throughout. Remarkably, the QBs slightly increase even as the noise strength rises to $\lambda=0.9$. The difference between the QB at $\lambda=0.9$ and the noiseless case is $\Delta=1.68\times 10^{-5}$ for F4 and $\Delta=3.73\times 10^{-5}$ for F9. This demonstrates that the optimal sensor networks remain robust under dephasing, with only minimal degradation in sensitivity even without re-optimization.
Figure 5: (a) QBs for nine qubits ($N=9$) using ansatze in Fig. 1(c); F9 gives the best QB, above the four-qubit case. (b) QBs versus $L_{1}$ for F9. (c) CBs versus  $L_{2}$ for F9.
@@
Conclusion. We proposed and demonstrated a network-based quantum sensing strategy for enhancing DM detection. By structuring superconducting qubits into networks with specific symmetries and optimizing their connections using VQM, we achieved substantial improvements in sensing precision compared to conventional GHZ-based protocols. Both the QB and CB were systematically minimized, and Bayesian estimation validated the enhanced sensitivity of the optimized networks. The optimal configurations maintained shallow circuit depths, making them feasible for implementation on NISQ devices. Our findings emphasize the significant impact of network design on quantum metrology performance and suggest that well-structured quantum network can serve as powerful platforms for DM searches and for a wide range of high-precision sensing applications. Future works may extend these results by exploring larger networks, adaptive strategies, and integration with error mitigation techniques to further push the limits of quantum-enhanced detection.
@@
Appendix: Variational Quantum Metrology Optimization Method. We detail the optimization process employed in variational quantum metrology (VQM). Consider the F4 configuration as an example, with the quantum circuit illustrated in Fig.  7. The system is initialized in the state $|\bm{0}\rangle$, which evolves to $|\Psi(\delta,\alpha,\bm{\theta})\rangle=V(\delta,\alpha)U(\bm{\theta})|\bm{0}\rangle.$ The QFI is given by
$Q=4\left(\langle\partial_{\delta}\Psi|\partial_{\delta}\Psi\rangle-\left|\langle\Psi|\partial_{\delta}\Psi\rangle\right|^{2}\right),$ (12)
where dependence on $\delta,\alpha,\bm{\theta}$ is omitted for brevity.
Consider the cost function as $\mathcal{C}(\bm{\theta})=1/Q$, and minimize it using the Adam optimizer. The update rule at iteration $t$ is
$\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon},$ (13)
where $m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})g_{t},\quad v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2},\hat{m}_{t}=\frac{m_{t}}{1-\beta_{1}^{t}},\quad\hat{v}_{t}=\frac{v_{t}}{1-\beta_{2}^{t}},$ and $g_{t}=\nabla_{\bm{\theta}}\mathcal{C}(\bm{\theta}_{t})$. We use $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. The process is repeated until convergence of $\mathcal{C}(\bm{\theta})$.
To evaluate the required derivatives, i.e., $|\partial_{\delta}\Psi\rangle$ and $\nabla_{\bm{\theta}}\mathcal{C}(\bm{\theta})$, we employ the parameter-shift rule [35].
Figure 6: QBs under dephasing noise for (a) F4 and (b) F9. Sensitivity decreases slightly with increasing $\lambda$, but even at $\lambda=0.9$ the rise is small: $\Delta=1.68\times 10^{-5}$ (a) and $\Delta=3.73\times 10^{-5}$ (b).
Figure 7: Quantum circuit for optimizing the QB in VQM.